OpenStudy (anonymous):

Write a rule for the linear function in the table. x f(x) –4 16 0 0 4 –16 8 –32 A. f(x) = –4x B. f(x) = x + 16 C. f(x) = 4x D. f(x) = x

OpenStudy (anonymous):

A. f(x) = –4x

OpenStudy (anonymous):

Ok. We're told that f(x) is linear, so it will be a polynomial of order 1. In other words, a line in the form of y=mx+b. Finding b is easy as we have f(0). Since f(0)=0, then 0=m0+b, which implies that b is 0. Then using any other set of points (say x=-4) we can find m:\[f(-4)=m(-4)+0=16\]\[\rightarrow m=-4\]So\[f(x)=-4x\]

OpenStudy (anonymous):

thank you